# Quantum graphs of homomorphisms

### v1

@@ Abstract
The finite quantum graphs in $\mathsf{qGph}$ are tracial, real, and self-adjoint, and the morphisms between them are CP morphisms that are adjoint to a unital *-homomorphism. We show that Weaver’s two notions of a CP morphism coincide in this context. We also show that every finite reflexive quantum graph is the confusability quantum graph of a quantum channel, answering a question of Daws.
@@ 1.2 The quantum graph of homomorphisms $G\to H$
Our approach adheres to orthodox noncommutative geometry. We generalize sets to discrete quantum spaces [37], i.e., quantum sets [23], and generalize relations to quantum relations [42], which correspond to projections in this setting [23, Theorems 3.6 and B.8]. Explicitly, in the meaning of the present article, a quantum graph is essentially a von Neumann algebra $\mathcal{M}\subseteq\mathcal{B}(H)$ and an ultraweakly closed operator system $\mathcal{R}\subseteq\mathcal{B}(H)$ such that $\mathcal{M}^{\prime}\cdot\mathcal{R}\cdot\mathcal{M}$$^{\prime}\subseteq\mathcal{R}$ and $\mathcal{R}^{\dagger}=\mathcal{R}$ (cf. [42, Definition 2.6(d)]) and such that $\mathcal{M}\cong\bigoplus_{i\in I}^{\ell^{\infty}}M_{n_{i}}(\mathbb{C})$ for some indexed family of positive integers $\{n_{i}\}_{i\in I}$, which may be infinite.
@@ 1.3 Comments
Context. The category $\mathsf{qGph}$ fits into prior literature in the following way: Modulo the choice of formalization, an object of $\mathsf{qGph}$ is a quantum graph in the sense of Weaver [41, Definition 2.6(d)] that need not be reflexive [41, Definition 2.4(d)(i)] and that is on a hereditarily atomic von Neumann algebra [23, Definition 5.3]. Therefore, the finite quantum graphs in $\mathsf{qGph}$ are the real, tracial, and self-adjoint (finite) quantum graphs in the terminology of [32].
@@
Acknowledgments. The first author was supported by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0041 and by the National Science Foundation under Award No. DMS-2231414. The second author was supported by the Austrian Science Fund (FWF) under Project DOI 10.55776/PAT6443523, by the Scientific Grant Agency of the Ministry of Education, Science, Research and Sport of the Slovak Republic and the Slovak Academy of Sciences (VEGA) under Grant No. 2/0128/24 and by the Slovak Research and Development Agency (APVV) under Grant No. APVV-22-0570.
@@ 2 Quantum sets
It is generally preferable to work with the von Neumann algebras $\ell^{\infty}(\mathcal{X})$ rather than the $C^{*}$-algebras $c_{0}(\mathcal{X})$ for two reasons. First, functions $\mathcal{X}\to\mathcal{Y}$ correspond to Woronowicz morphisms $c_{0}(\mathcal{Y})\to c_{0}(\mathcal{X})$ and to unital normal $\dagger$-homomorphisms $\ell^{\infty}(\mathcal{Y})\to\ell^{\infty}(\mathcal{X})$, and the latter notion is technically simpler. Second, relations $\mathcal{X}\to\mathcal{Y}$ correspond to quantum relations [41] between $\ell^{\infty}(\mathcal{X})$ and $\ell^{\infty}(\mathcal{Y})$, and while we speculate that they correspond to a class of $C^{*}$-correspondences [36] between $c_{0}(\mathcal{X})$ and $c_{0}(\mathcal{Y})$ too, we are not aware of this connection in the literature. Up to isomorphism, a von Neumann algebra is of the form $\ell^{\infty}(\mathcal{X})$ iff it is hereditarily atomic [43, Proposition 5.4].
A relation $R$ from $\mathcal{X}$ to $\mathcal{Y}$ is a choice of subspaces $R(X,Y)\leq L(X,Y)$ for all $X\in\mathrm{At}(\mathcal{X})$ and $Y\in\mathrm{At}(\mathcal{Y})$. Quantum sets and relations form a dagger compact category $\mathsf{qRel}$ in a straightforward way [43, Theorem. 3.6], which may be guessed by the reader. A dagger compact category, i.e., a strongly compact category [1], is a symmetric monoidal category that is further equipped with a choice of dual objects and a compatible contravariant functor $(-)^{\dagger}$. The category of $\mathsf{FinHilb}$ of finite-dimensional Hilbert spaces and linear operators and the category $\mathsf{Rel}$ of sets and relations are two prominent examples.
@@
A function $F\colon\mathcal{X}\to\mathcal{Y}$ is a relation from $\mathcal{X}$ to $\mathcal{Y}$ such that $F^{\dagger}\circ F\geq\mathrm{Id}_{\mathcal{X}}$ and $F\circ F^{\dagger}\leq\mathrm{Id}_{\mathcal{Y}}$, where $\mathrm{Id}_{\mathcal{X}}$ and $\mathrm{Id}_{\mathcal{Y}}$ are identity relations. The resulting category $\mathsf{qSet}$ of quantum sets and functions is dual to the category of hereditarily atomic von Neumann algebras and unital normal $\dagger$-homomorphisms [43, Theorem. 7.4].
@@ 6 Homomorphisms as quantum operations
In this section, we identify homomorphisms between finite quantum graphs with certain quantum operations, confirming that Definition˜3.1 is, for finite quantum graphs, a special case of [42, Definition 8.5]. The latter definition is a quantum generalization of substochastic graph homomorphisms, whereas Definition˜3.1 is a quantum generalization of graph homomorphisms in the standard sense.
@@
Claims 6 and 7 follow immediately from [42, Theorem 8.4] and [42, Theorem 9.2], respectively. ∎
@@
    $\varphi$ is a CP morphism [42, Definition 8.5] from $(\mathcal{M},R)$ to $(\mathcal{N},S)$,
@@
As Weaver observes in [42, section 9], the condition $\mathcal{R}\subseteq\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{S}\cdot\mathcal{T}_{\varphi}$ defines an alternative notion of a CP morphism, which is strictly weaker when $\varphi$ is trace-preserving. We now prove that when $\varphi\colon\mathcal{M}\to\mathcal{N}$ is a trace-preserving $\dagger$-cohomomorphism, these two conditions are equivalent and define exactly the graph homomorphisms of Definition˜3.1.
@@
When $\mathcal{M}\subseteq M_{m}(\mathbb{C})\cong L(\mathbb{C}^{m})$ is in standard form [17], $m=\dim\mathcal{M}$, and the trace on $M_{m}(\mathbb{C})$ restricts to the adjusted trace [25, Remark 1.4] or Markov trace [10, section 7] on $\mathcal{M}$, which normalizes to the unique tracial $\delta$-form on $\mathcal{M}$ [3]. In this case, $\mu\colon\mathcal{M}\to\mathcal{M}\otimes\mathcal{M}$ is the comultiplication of $\mathcal{M}$ as a special unitary $\dagger$-Frobenius involution monoid [40, Theorem 4.6].
To establish a connection between Weaver’s CP morphisms [42, Definition 8.5] and our homomorphisms of quantum graphs (Definition˜3.1), we show that the assignment $\varphi\mapsto\mathcal{T}_{\varphi}$ in Item˜6.1(2) is a one-to-one correspondence between trace-preserving $\dagger$-cohomomorphisms and quantum functions [21].
@@
We conclude the article by proving that, in the context of finite quantum graphs, our homomorphisms $G\to H$ correspond to those trace-preserving $\dagger$-cohomomorphisms $\ell(\mathcal{V}_{G})\to\ell(\mathcal{V}_{H})$ that are CP morphisms in either of Weaver’s senses [42, section 9], which become equivalent.
@@
A trace-preserving $\dagger$-homomorphism $\mathcal{M}\to\mathcal{N}$ is a quantum channel in the sense of being a trace-preserving completely positive map. In order to provide an interpretation of the morphisms in Theorem˜6.6 within quantum information theory, we verify that every reflexive, symmetric quantum relation $\mathcal{R}$ on $\mathcal{M}$ is the confusability quantum graph of a quantum channel (Item˜6.1(4)), answering a question of Daws [10, section 6.2].
@@
Let $\mathcal{M}\subseteq M_{m}(\mathbb{C})$ be a $\dagger$-matrix algebra, and let $\mathcal{R}\subseteq M_{m}(\mathbb{C})$ be a reflexive, symmetric quantum relation on $\mathcal{M}$ [41, Definition 2.4(d)]. Then, there exist a positive integer $n$ and a trace-preserving completely positive map $\varphi\colon\mathcal{M}\to M_{n}(\mathbb{C})$ such that $\mathcal{R}=\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{T}_{\varphi}$.
@@
In the case $\mathcal{M}=M_{m}(\mathbb{C})$, this proposition is widely known [12, Lemma 2]. Thus, we obtain a trace-preserving completely positive map $\psi\colon M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ such that $\mathcal{T}_{\psi}^{\dagger}\cdot\mathcal{T}_{\psi}=\mathcal{R}$. Let $\iota\colon\mathcal{M}\to M_{m}(\mathbb{C})$ be the inclusion map, which is also trace-preserving and completely positive. The Kraus decomposition of $\iota$ is given by $\iota(a)=1_{m}a1_{m}$, so $\mathcal{T}_{\iota}=\mathcal{M}^{\prime}$. Let $\varphi=\psi\circ\iota$. The functoriality of the assignment $\varphi\mapsto\mathcal{T}_{\varphi}$ implies that $\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{T}_{\varphi}=\mathcal{T}_{\iota}^{\dagger}\cdot\mathcal{T}_{\psi}^{\dagger}\cdot\mathcal{T}_{\psi}\cdot T_{\iota}=\mathcal{M}^{\prime}\cdot\mathcal{R}\cdot\mathcal{M}^{\prime}=\mathcal{R}.$
@@
    trace-preserving $\dagger$-cohomomorphisms $\varphi\colon\ell(\mathcal{V}_{G})\to\ell(\mathcal{V}_{H})$ that are CP morphisms for the quantum relations $\mathcal{R}_{G}$ and $\mathcal{R}_{H}$ [42, Definition 8.5].

### v2

@@ Abstract
The finite quantum graphs in $\mathsf{qGph}$ are tracial, real, and self-adjoint, and the morphisms between them are CP morphisms that are adjoint to a unital *-homomorphism. We prove that Weaver’s two notions of a CP morphism coincide in this context. We also include a short proof that every finite reflexive quantum graph is the confusability quantum graph of a quantum channel.
@@ 1.2 The quantum graph of homomorphisms $G\to H$
Our approach adheres to orthodox noncommutative geometry. We generalize sets to discrete quantum spaces [37], i.e., quantum sets [23], and generalize relations to quantum relations [43], which correspond to projections in this setting [23, Theorems 3.6 and B.8]. Explicitly, in the meaning of the present article, a quantum graph is essentially a von Neumann algebra $\mathcal{M}\subseteq\mathcal{B}(H)$ and an ultraweakly closed operator system $\mathcal{R}\subseteq\mathcal{B}(H)$ such that $\mathcal{M}^{\prime}\cdot\mathcal{R}\cdot\mathcal{M}$$^{\prime}\subseteq\mathcal{R}$ and $\mathcal{R}^{\dagger}=\mathcal{R}$ (cf. [43, Definition 2.6(d)]) and such that $\mathcal{M}\cong\bigoplus_{i\in I}^{\ell^{\infty}}M_{n_{i}}(\mathbb{C})$ for some indexed family of positive integers $\{n_{i}\}_{i\in I}$, which may be infinite.
@@ 1.3 Comments
Context. The category $\mathsf{qGph}$ fits into prior literature in the following way: Modulo the choice of formalization, an object of $\mathsf{qGph}$ is a quantum graph in the sense of Weaver [42, Definition 2.6(d)] that need not be reflexive [42, Definition 2.4(d)(i)] and that is on a hereditarily atomic von Neumann algebra [23, Definition 5.3]. Therefore, the finite quantum graphs in $\mathsf{qGph}$ are the real, tracial, and self-adjoint (finite) quantum graphs in the terminology of [32].
@@
Acknowledgments. The first author was supported by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0041 and by the National Science Foundation under Award No. DMS-2231414. The second author was supported by the Austrian Science Fund (FWF) under Project DOI 10.55776/ PAT6443523.
@@ 2 Quantum sets
It is generally preferable to work with the von Neumann algebras $\ell^{\infty}(\mathcal{X})$ rather than the $C^{*}$-algebras $c_{0}(\mathcal{X})$ for two reasons. First, functions $\mathcal{X}\to\mathcal{Y}$ correspond to Woronowicz morphisms $c_{0}(\mathcal{Y})\to c_{0}(\mathcal{X})$ and to unital normal $\dagger$-homomorphisms $\ell^{\infty}(\mathcal{Y})\to\ell^{\infty}(\mathcal{X})$, and the latter notion is technically simpler. Second, relations $\mathcal{X}\to\mathcal{Y}$ correspond to quantum relations [42] between $\ell^{\infty}(\mathcal{X})$ and $\ell^{\infty}(\mathcal{Y})$, and while we speculate that they correspond to a class of $C^{*}$-correspondences [36] between $c_{0}(\mathcal{X})$ and $c_{0}(\mathcal{Y})$ too, we are not aware of this connection in the literature. Up to isomorphism, a von Neumann algebra is of the form $\ell^{\infty}(\mathcal{X})$ iff it is hereditarily atomic [44, Proposition 5.4].
A relation $R$ from $\mathcal{X}$ to $\mathcal{Y}$ is a choice of subspaces $R(X,Y)\leq L(X,Y)$ for all $X\in\mathrm{At}(\mathcal{X})$ and $Y\in\mathrm{At}(\mathcal{Y})$. Quantum sets and relations form a dagger compact category $\mathsf{qRel}$ in a straightforward way [44, Theorem. 3.6], which may be guessed by the reader. A dagger compact category, i.e., a strongly compact category [1], is a symmetric monoidal category that is further equipped with a choice of dual objects and a compatible contravariant functor $(-)^{\dagger}$. The category of $\mathsf{FinHilb}$ of finite-dimensional Hilbert spaces and linear operators and the category $\mathsf{Rel}$ of sets and relations are two prominent examples.
@@
A function $F\colon\mathcal{X}\to\mathcal{Y}$ is a relation from $\mathcal{X}$ to $\mathcal{Y}$ such that $F^{\dagger}\circ F\geq\mathrm{Id}_{\mathcal{X}}$ and $F\circ F^{\dagger}\leq\mathrm{Id}_{\mathcal{Y}}$, where $\mathrm{Id}_{\mathcal{X}}$ and $\mathrm{Id}_{\mathcal{Y}}$ are identity relations. The resulting category $\mathsf{qSet}$ of quantum sets and functions is dual to the category of hereditarily atomic von Neumann algebras and unital normal $\dagger$-homomorphisms [44, Theorem. 7.4].
@@ 6 Homomorphisms as quantum operations
In this section, we identify homomorphisms between finite quantum graphs with certain quantum operations, confirming that Definition˜3.1 is, for finite quantum graphs, a special case of [43, Definition 8.5]. The latter definition is a quantum generalization of substochastic graph homomorphisms, whereas Definition˜3.1 is a quantum generalization of graph homomorphisms in the standard sense.
@@
Claims 6 and 7 follow immediately from [43, Theorem 8.4] and [43, Theorem 9.2], respectively. ∎
@@
    $\varphi$ is a CP morphism [43, Definition 8.5] from $(\mathcal{M},R)$ to $(\mathcal{N},S)$,
@@
As Weaver observes in [43, section 9], the condition $\mathcal{R}\subseteq\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{S}\cdot\mathcal{T}_{\varphi}$ defines an alternative notion of a CP morphism, which is strictly weaker when $\varphi$ is trace-preserving. We now prove that when $\varphi\colon\mathcal{M}\to\mathcal{N}$ is a trace-preserving $\dagger$-cohomomorphism, these two conditions are equivalent and define exactly the graph homomorphisms of Definition˜3.1.
@@
When $\mathcal{M}\subseteq M_{m}(\mathbb{C})\cong L(\mathbb{C}^{m})$ is in standard form [17], $m=\dim\mathcal{M}$, and the trace on $M_{m}(\mathbb{C})$ restricts to the adjusted trace [25, Remark 1.4] or Markov trace [10, section 7] on $\mathcal{M}$, which normalizes to the unique tracial $\delta$-form on $\mathcal{M}$ [3]. In this case, $\mu\colon\mathcal{M}\to\mathcal{M}\otimes\mathcal{M}$ is the comultiplication of $\mathcal{M}$ as a special unitary $\dagger$-Frobenius involution monoid [41, Theorem 4.6].
To establish a connection between Weaver’s CP morphisms [43, Definition 8.5] and our homomorphisms of quantum graphs (Definition˜3.1), we show that the assignment $\varphi\mapsto\mathcal{T}_{\varphi}$ in Item˜6.1(2) is a one-to-one correspondence between trace-preserving $\dagger$-cohomomorphisms and quantum functions [21].
@@
We conclude the article by proving that, in the context of finite quantum graphs, our homomorphisms $G\to H$ correspond to those trace-preserving $\dagger$-cohomomorphisms $\ell(\mathcal{V}_{G})\to\ell(\mathcal{V}_{H})$ that are CP morphisms in either of Weaver’s senses [43, section 9], which become equivalent.
@@
A trace-preserving $\dagger$-homomorphism $\mathcal{M}\to\mathcal{N}$ is a quantum channel in the sense of being a trace-preserving completely positive map. In order to provide an interpretation of the morphisms in Theorem˜6.6 within quantum information theory, we appeal to the fact that every reflexive, symmetric quantum relation $\mathcal{R}$ on $\mathcal{M}$ is the confusability quantum graph of a quantum channel in the sense of Item˜6.1(4). This was established in greater generality by Verdon [40, Proposition 3.12], answering a question of Daws [10, section 6.2]. We provide a short proof of this fact.
@@
Let $\mathcal{M}\subseteq M_{m}(\mathbb{C})$ be a $\dagger$-matrix algebra, and let $\mathcal{R}\subseteq M_{m}(\mathbb{C})$ be a reflexive, symmetric quantum relation on $\mathcal{M}$ [42, Definition 2.4(d)]. Then, there exist a positive integer $n$ and a trace-preserving completely positive map $\varphi\colon\mathcal{M}\to M_{n}(\mathbb{C})$ such that $\mathcal{R}=\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{T}_{\varphi}$.
@@
In the case $\mathcal{M}=M_{m}(\mathbb{C})$, this proposition is widely known [12, Lemma 2]. Thus, we obtain a trace-preserving completely positive map $\psi\colon M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ such that $\mathcal{T}_{\psi}^{\dagger}\cdot\mathcal{T}_{\psi}=\mathcal{R}$. Let $\iota\colon\mathcal{M}\to M_{m}(\mathbb{C})$ be the inclusion map, which is also trace-preserving and completely positive. The Kraus decomposition of $\iota$ is given by $\iota(a)=1_{m}a1_{m}$, so $\mathcal{T}_{\iota}=\mathcal{M}^{\prime}$. Let $\varphi=\psi\circ\iota$. The functoriality of the assignment $\varphi\mapsto\mathcal{T}_{\varphi}$ implies that $\mathcal{T}_{\varphi}^{\dagger}\cdot\mathcal{T}_{\varphi}=\mathcal{T}_{\iota}^{\dagger}\cdot\mathcal{T}_{\psi}^{\dagger}\cdot\mathcal{T}_{\psi}\cdot\mathcal{T}_{\iota}=\mathcal{M}^{\prime}\cdot\mathcal{R}\cdot\mathcal{M}^{\prime}=\mathcal{R}.$
@@
    trace-preserving $\dagger$-cohomomorphisms $\varphi\colon\ell(\mathcal{V}_{G})\to\ell(\mathcal{V}_{H})$ that are CP morphisms for the quantum relations $\mathcal{R}_{G}$ and $\mathcal{R}_{H}$ [43, Definition 8.5].